# Boundary Constraint-free Biomechanical Model-Based Surface Matching for Intraoperative Liver Deformation Correction

### v1

@@ V-E Parameter Sensitivity Analysis
The proposed method relies on three parameters: (1) $k_{ss}$, a stabilizing parameter for the solution; (2) $\beta$, a parameter regulating the spatial smoothness of the estimated forces; and (3) Poisson’s ratio $\nu$, defining the compressibility of the tissue. To understand the sensitivity of our method to the parameters, we vary each parameter individually while keeping the others constant. We examine their impact on registration errors using the in silico phantom and the partial case of No.1 phantom in Sec. V-B.
Figure 8: Parameter sensitivity analysis for $k_{ss}$, $\beta$, and $\nu$ performed on the in silico phantom (first column) and the partial case of Phantom No. 1 (second column). The median and mean values are represented by orange lines and crosses, respectively.
As illustrated in Fig. 8, the registration outcomes exhibit a degree of sensitivity to the parameter $k_{ss}$. The default setting of $k_{ss}=0.01$ shows competitive performance, outperforming other values on the in silico phantom. However, the other value $k_{ss}=0.001$ yields the minimum errors on the phantom data. In previous sections, we found that $k_{ss}=0.01$ provides reasonable errors across all testing samples. Hence, while refining this parameter could potentially enhance results, $k_{ss}$ values between 0.01 and 0.1 consistently yield satisfactory outcomes. To be noted, when $k_{ss}=0$ (no added soft spring stiffness), the registration errors are all obviously larger than the errors when we set $k_{ss}$ values between 0.01 and 0.1.
For the in silico phantom without added noise, increasing $\beta$ leads to escalating errors. However, for dealing with phantoms that include noise, such as segmentation noise in CT scans, varying $\beta$ results in stable performance.
The default setting of Poisson’s ratio $\nu=0.49$ corresponds to nearly incompressible tissue behavior. While some prior studies [4, 1] adopt $\nu=0.45$, our experiments involving variations from 0.45 to 0.49 show stable performance on both the in silico and in vitro phantoms.
@@ VI Discussion
We have presented a novel non-rigid 3D-3D registration method that incorporates FEM within the data term and eliminates the need to identify zero boundary conditions and force locations.
@@
To eliminate the need for priors, we first introduce a soft spring modification to the stiffness matrix. This modification has been demonstrated to substantially reduce registration errors, as evidenced in the first row of Fig 8. Secondly, the method is designed to search for optimal forces autonomously. In the absence of manually identified ZBC and force locations, this strategy is more beneficial than the previous approaches [4] of constraining force locations to predefined areas, such as the posterior surface, shown in Fig. 4 (b). Furthermore, our method incorporates L2 regularization that imposes spatial smoothness. The L2 regularization ensures the model remains robust under conditions where the intraoperative point cloud is noisy, as shown in Fig. 4 (d). In addition, our proposed optimal step determination, as demonstrated in Fig. 4 (c), eliminates the need for manual tuning of the step size parameter and leads to quicker convergence and lower error.
@@
We have evaluated our method across various datasets, encompassing our phantom datasets, the Open-CAS dataset, and the Sparse dataset. Moreover, in direct comparisons with two alternative methods, V2Snet and GMM-FEM, our approach consistently outperforms or shows comparable performance, as elaborated in Sections V-B, V-C, and V-D.
@@
While GMM-FEM incorporates FEM into the strain energy, regularizing the deformation field estimated by the CPD registration method, it does not guarantee faithful preservation of the original geometry (i.e., vertex connectivity) of the deformed surface. This becomes apparent in situations with noisy intraoperative point cloud data, as illustrated in Fig. 6. In contrast, our approach, where FEM inherently guides the deformation, ensures the maintenance of volumetric vertex connectivity, preserving the geometry during deformation. Furthermore, as demonstrated in Section V-B, GMM-FEM exhibits sensitivity to the visibility of the intraoperative surface, as highlighted in its original paper [17].
V2Snet utilizes FEM to construct a training dataset [20]. While it shows competitive results in specific instances, its overall performance and generalization capabilities require improvement. Nevertheless, the use of learning-based methods, offering prior information, has the potential to assist optimization methods in achieving faster and more accurate results.
@@
To be noted, before performing the non-rigid registration, it is crucial to (1) perform a optimal rigid alignment, (2) minimize the intraoperative liver surface reconstruction errors, and (3) ensure a high quality surface characterized by sufficient convergence. These considerations are prerequisites for the problem. Otherwise, our method, like any other, may result in inaccurate registration. However, such registration inaccuracies may be mitigated, provided other constraints besides the intraoperative surface are available, such as vascular features and ligaments from ultrasound [1].
@@ VII Conclusion
We have presented a novel non-rigid 3D-3D registration method that seamlessly integrates a biomechanical model into the surface-matching term without relying on priors. By embedding the biomechanical model in the surface-matching term, the estimated deformation is forced to preserve the geometry of the preoperative model. Experiments on our in silico phantom and in vitro phantom datasets and two public datasets demonstrate the method’s efficacy in achieving accurate registration results and preserving original geometry. Furthermore, our constructed phantom dataset and method move the benchmarking liver registration, a valuable addition considering the limited availability of open-source datasets and methods. In future work, we plan to implement the algorithm on GPUs, research the integration of learning-based approaches.

### v2

@@ V-E Parameter Sensitivity Analysis
The proposed method relies on three parameters: (1) $k_{ss}$, a stabilizing parameter for the solution; (2) Poisson’s ratio $\nu$, defining the compressibility of the tissue; and (3) iteration number. To understand the sensitivity of our method to these parameters, we vary each parameter individually while keeping the others constant. We examined their impact on registration error and conducted statistical Kolmogorov–Smirnov tests at a significance level of $\alpha=0.05$, using the entire sparse dataset. The results are shown in Fig. 8.
The registration exhibit a degree of sensitivity to the parameter $k_{ss}$. The default setting of $k_{ss}=0.01$ outperforms other settings ($p<0.001$). The proposed method is not sensitive to Poisson’s ratio $\nu$; hence, no statistical significance is found. The default setting of Poisson’s ratio $\nu=0.49$ corresponds to nearly incompressible tissue behavior. While some prior studies [4, 1] adopt $\nu=0.45$, our experiments involving variations from 0.45 to 0.49 show stable performance. Lastly, when we reduce the iteration number to 50, statistical significance is found ($p<0.001$), suggesting the default iteration number 200 could be further tuned to reduce computation.
@@ VI Discussion
VI-A Summary of Main Findings
@@
We first introduce a soft spring modification to the stiffness matrix to eliminate the need for priors. This modification has been demonstrated to substantially reduce registration errors, as evidenced in Fig 8. Secondly, the method is designed to search for optimal forces autonomously. In the absence of manually identified ZBC and force locations, this strategy is more beneficial than the previous approaches [4] of constraining force locations to predefined areas, such as the posterior surface, shown in Fig. 5 (b). In addition, our proposed optimal step determination, as demonstrated in Fig. 5 (c), eliminates the need for manual tuning of the step size parameter and leads to quicker convergence and lower error.
@@
We have evaluated our method across various datasets, encompassing our phantom datasets, the Open-CAS dataset, and the Sparse dataset. Moreover, in comparisons with two alternative open source methods (V2Snet and GMM-FEM) and closed source methods reported in the Sparse Data Non-rigid Registration Challenge [38, 22], our approach consistently outperforms or shows comparable performance, as elaborated in Sections V-B, V-C, and V-D.
@@
In comparison with open source methods, while GMM-FEM incorporates a FEM into the strain energy, regularizing the deformation field estimated by the CPD registration method, it does not guarantee faithful preservation of the original geometry (i.e., vertex connectivity) of the deformed surface. This becomes apparent in situations with noisy intraoperative point cloud data, as illustrated in Fig. 7. In contrast, our approach, where FEM inherently guides the deformation, ensures the maintenance of volumetric vertex connectivity, preserving the geometry during deformation. Furthermore, as demonstrated in Section V-B, GMM-FEM exhibits sensitivity to the visibility of the intraoperative surface, as highlighted in its original paper [17]. V2Snet utilizes FEM to construct a training dataset [20]. While it shows competitive results in specific instances, its overall performance and generalization capabilities require improvement. Nevertheless, the use of learning-based methods, offering prior information, has the potential to assist optimization methods in achieving faster and more accurate results.
The sparse dataset enables comparison of our proposed method with other closed-source methods, particularly regarding surface visibility, initial alignment, and noise magnitude. Our method demonstrates robustness to these factors within reasonable variations and consistently achieves the lowest registration errors, as shown in Table III, IV, and V.
The proposed method is straightforward and effective, in contrast to registration methods driven by both surface-matching and strain energy terms. Our approach relies solely on the surface-matching term embedded within the FEM, eliminating the need to iteratively determine boundary conditions [1, 4] or manually specify them [16]. Additionally, our method depends on only a few parameters, making it easy to adapt to other datasets. Section V-E provides a sensitivity analysis of these parameters. The stabilizing parameter $k_{ss}$ should be set carefully, though we recommend against extensive tuning, as the default parameter value performs well across the samples included in this study. Potential future work includes automating the process of determining the $k_{ss}$ value.
VI-B Linear Elastic Model and Deformation
We have shown that our algorithm’s linear elastic model version is parameter robust and leads to accurate registrations for datasets with deformations typically observed in laparoscopic [9] and open liver surgery [23]. The linear elastic model is generally employed by most state-of-the-art methods [16, 1, 38]. In fact, all biomechanical-based methods reported in the Sparse Challenge dataset adopted the linear elastic model. The maximum TRE reported in the sparse data is 11.9 $\pm$ 3.7, consistent with clinically observed deformation magnitudes for open liver surgery reported in Ref. [23, 39].
Liver deformation in laparoscopic interventions can be larger due to the creation of pneumoperitoneum. The Open-CAS dataset features even more significant deformations than the Sparse Challenge dataset. Even following the initial rigid alignment correction, mean TRE still ranges from 5.88 mm to 13.95 mm, which reflects the typical deformations observed in laparoscopic liver surgeries reported in [9], where surface deformations ranged from 7.7 mm to 12.6 mm at 14 mmHg and from 6.2 mm to 11.9 mm at 7 mmHg. Our proposed registration method achieved mean TRE values below 5 mm across all these samples despite featuring large deformations, with the exception of one sample where the intraoperative surface was highly compromised due to a highly inaccurate stereo reconstruction. Hence, the widely adopted linear elastic biomechanical approach can yield reasonable registration accuracy, even within these typically large deformation ranges. This observation is also shared by previous works [4, 39, 45] that also suggested that a linear elastic model is perhaps adequate to achieve sufficiently accurate non-rigid organ registration.
The liver can also be represented using a nonlinear biomechanical model. While nonlinear models may be considered superior to linear elastic models at capturing large deformations, the material properties required for such models to be sufficiently realistic and accurate vary significantly among patients and are difficult to measure, especially during surgery. Additionally, using a nonlinear model increases computational complexity, as the tangent stiffness matrix must be recalculated with each algorithm iteration. To incorporate a nonlinear model into our method, line 5 of Algorithm 1 can be modified to utilize the FEM for calculating deformations and the tangent stiffness matrix. Various nonlinear models can be selected for this purpose. For further details, we refer the reader to the works described in [45, 15, 3, 46], leaving this aspect for future exploration.
VI-C Potential Challenges
The intraoperative point cloud, collected using an optically tracked stylus, is very sparse compared to other imaging modalities. However, despite this sparsity, the proposed method still achieved accurate results. To address extremely sparse intraoperative point clouds, a grid-fitting interpolation method similar to the one used by Collins et al. [37] could be employed to recover denser surface points. However, it is recommended to avoid such sparse conditions during data collection, and hence, an extremely sparse data analysis would become essentially unrealistic.
@@ VI-C Potential Challenges
Before performing non-rigid registration, it is essential to (1) minimize intraoperative liver surface reconstruction errors, (2) ensure a high-quality surface with sufficient convergence, and (3) achieve accurate rigid alignment. These steps are critical prerequisites; otherwise, our method, like others, may yield inaccurate results. Integrating real-time imaging modalities, such as ultrasound or cone-beam CT, could potentially relax these requirements by providing additional constraints, such as vascular features (i.e., ligaments visible in ultrasound) or a complete intraoperative surface from cone-beam CT.
@@ VII Conclusion
We have presented a unique non-rigid 3D-3D registration method that seamlessly integrates a biomechanical model into the surface-matching term without relying on priors. By embedding the biomechanical model in the surface-matching term, the estimated deformation is forced to preserve the geometry of the preoperative model without an additional strain energy term. Experiments on our in silico phantom, in vitro phantom datasets, and two publicly available datasets demonstrate the method’s efficacy in achieving accurate registration results and preserving original geometry. Furthermore, our constructed phantom dataset and registration method represent significant contributions to liver registration benchmarking and, hence, valuable additions considering the limited availability of open-source datasets and methods.
As part of our future work, we intend to research the effects of nonlinear hyperelastic materials, incorporating learning-based methods, automating the process of determining the stabilizing parameter and the potential combination of this proposed method with other real-time intraoperative imaging modalities. We also note that the 3D-3D non-rigid registration problem tackled here shares many similarities with those used for other organs, such as the prostate, kidney, brain, and breast, and hence, we would be interested in investigating the performance of the proposed method on other organs.